OpenStudy (anonymous):

Write the equation of the line passing through the points (-6,1) and (-4,2).

OpenStudy (anonymous):

Find the slope by using the formula: (y2-y1)/(x2-x1) = m then use point slope form: y-y1=m(x-x1)

OpenStudy (anonymous):

what does that mean?

OpenStudy (anonymous):

the slope is -1?

OpenStudy (anonymous):

(x1,y1) (x2,y2) (-6, 1) (-4, 2) now just plug it in

OpenStudy (anonymous):

I did that part how do i do the other equation?

OpenStudy (anonymous):

(2-1)/(-4-(-6)) (1)/(2) != -1

OpenStudy (anonymous):

it's 1/2. that is your SLOPE now use the point slope form. plug in M, x1,y1 then set it equal to y. u should have y=mx+b